2 identical spheres are place at a distance so that force of attraction between them is s if 50% of mass of 1 sphere is transferred to other sphere then what will be the change of force of attraction


1
Expert's answer
2020-08-06T16:44:53-0400

Let's denote masses of each sphere as "m". At the initial distance "r" the force of attraction is "F_1". According to the Newton's gravity law, this force will be:

"F_1 = G\dfrac{m\cdot m}{r^2} = G\dfrac{m^2}{r^2}"

After the redistribution of masses, the first sphere will have the mass "m\/2" and the second one - "3m\/2". Then the attraction force will be:


"F_2 = G\dfrac{\frac{m}{2}\cdot \frac{3m}{2}}{r^2} = \dfrac{3}{4}G\dfrac{m^2}{r^2}"

Dividing the second force by the first, obtain:


"\dfrac{F_2}{F_1} = \dfrac34"

Thus, the force of attraction becomes 3/4 times less, then the initial one.


Answer. 3/4 times less.